OpenStudy (bookworm14):

Find the area of a regular pentagon with a radius of 5m. I got 43 is that correct?

OpenStudy (anonymous):

how did you get 43

OpenStudy (bookworm14):

wait , lemme rework this real quick, i think i messed up

OpenStudy (bookworm14):

i reworked it and got 59.3 is that correct?

OpenStudy (anonymous):

please wait let me solve mine real quick

OpenStudy (bookworm14):

ok :)

OpenStudy (anonymous):

I think that should help u sorry my calculator is messing up right now

OpenStudy (bookworm14):

im watching it now :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It should probably help u

OpenStudy (anonymous):

Please tell me if it helped

OpenStudy (whpalmer4):

I get 59.441, not 59.3. Are you evaluating \[\frac{125}{2}\sqrt{\frac{5}{8} + \frac{\sqrt{5}}8}\]

OpenStudy (bookworm14):

I am doing 1/2 * apothem * perimeter with values: 1/2 * 29.38 * 4.04 = 59.34 and i have to round the the nearest tenth (according to my directions) and that for me equals 59.3

OpenStudy (bookworm14):

@whpalmer4

OpenStudy (whpalmer4):

@bookworm14 you are using inaccurate intermediate values, which is distorting the final answer. You truncated \[5\sin(36^\circ)\approx 29.389\]to \(29.38\) instead of rounding it to \(29.39\) and you did the same with the other side. The correct area is \(\approx 59.441032268\), which rounded to the nearest tenth will give a different answer than yours. Instead of finding the length of one side of the triangle, writing down the first few digits, finding the second side, writing down a few digits, then punching those truncated or even rounded values back in, do the whole calculation in one shot: \[10*(\frac{!}2*(5\sin(36^\circ)*5\cos(36^\circ))\approx 59.441032268\]